OpenStudy (anonymous):

Which of the following points are NOT on the graph of y = 2x?

OpenStudy (across):

This is not a well-posed question.

OpenStudy (unklerhaukus):

\[y(x)=2x\qquad\qquad (2x,y)\] \[y(1)=2\times 1=2\qquad (1,2)\]\[y(2)=2\times 2=4\qquad(2,4)\]\[y(3)=2\times 3=6\qquad (3,6)\]